OpenStudy (anonymous):

determine convergence or divergence using appropriate test, identify test. infnity over E over n=2 (((-1)^n)/(n ln n))

OpenStudy (amistre64):

\[\sum_{n=2}^{\infty} f(n)\]

OpenStudy (anonymous):

\[\sum_{n=2}^{\infty}((-1)^{n}/n \times \ln(n))\] i suppose, converges with leibnits crit

OpenStudy (anonymous):

leibnitz*

OpenStudy (anonymous):

i do not know this

OpenStudy (amistre64):

\[\frac{(-1)^n}{n\ ln (n)}\]it flips back and forth from - to + right?

OpenStudy (amistre64):

but does it converge between them ... hmmm

OpenStudy (anonymous):

sure terms go to zero and it alternates

OpenStudy (amistre64):

ln(inf) is a large number; so yes, i agree. it goes to zero while alternating

OpenStudy (anonymous):

large number

OpenStudy (anonymous):

i get it is an alternating series , but what test would i use to tell if conv or diverging

OpenStudy (amistre64):

can we |restrain it| inside an absolute value?

OpenStudy (amistre64):

doesnt converge with absolutes

OpenStudy (amistre64):

can we try integral test?

OpenStudy (anonymous):

integral can be done, but is unnecesary

OpenStudy (amistre64):

i cant recall too many more tests; ratio test? comparison to familiar stuff tests?

OpenStudy (amistre64):

\[\sum\frac{-1^n}{n\ \sqrt{n}}\] converges

OpenStudy (anonymous):

this one is not absolutely convergent.

OpenStudy (anonymous):

ratio,root tests

OpenStudy (amistre64):

right, since the absolutes diverge :)

OpenStudy (anonymous):

for alternating series the leibnitz test is used, , it states that for \[\sum_{a}^{\infty}(f(x))\] if \[\lim_{x \rightarrow \infty}(f(x))\] converges, then (-1)^n*f(x) also converges

OpenStudy (anonymous):

made a little mistake, a (-1)^n and n=a should be written insidde the sum

OpenStudy (anonymous):

thank you everyone